I am having problems figuring out where to even start with the following, even just a hit would be appreciated.
ABC Car Rental Company determined that the distance traveled per car per year is normally
distributed, with a standard deviation of 645 miles. If 12% of the cars travel more than
55,000 miles, what is the mean distance traveled per year?

Answer by ewatrrr(24785)

88% ≤ 55,000
X the mean distance traveled per year
z = invNorm(.88) = (55,000 - X)/645
645invNorm(.88) = 55,000 - X
X = 55,000 - 645invNorm(.88)
X = 55,000 - 645(1.175) = 54,242.125, the mean distance traveled per year